OpenStudy (ccchristian1):

A school fair ticket costs $8 per adult and $1 per child. On a certain day, the total number of adults (a) and children (c) who went to the fair was 30, and the total money collected was $100. Which of the following options represents the number of children and the number of adults who attended the fair that day, and the pair of equations that can be solved to find the numbers? giving medals

OpenStudy (ccchristian1):

answers options : 1. 20 children and 10 adults Equation 1: a + c = 30 Equation 2: 8a + c = 100 2. 10 children and 20 adults Equation 1: a + c = 30 Equation 2: 8a − c = 100 3. 10 children and 20 adults Equation 1: a + c = 30 Equation 2: 8a + c = 100 4. 20 children and 10 adults Equation 1: a + c = 30 Equation 2: 8a − c = 100

OpenStudy (ccchristian1):

@TheWeeknd16 whats the answer

OpenStudy (theweeknd16):

You just need to find the correct equation ?

OpenStudy (ccchristian1):

will it be A or D? @TheWeeknd16

OpenStudy (theweeknd16):

im not sure. Try putting the question on www.jishka.com

OpenStudy (ccchristian1):

@HalpPlee

OpenStudy (ccchristian1):

@Savage16

OpenStudy (ccchristian1):

@triciaal

OpenStudy (ccchristian1):

@nincompoop

OpenStudy (theweeknd16):

@dan815 @too_smart143 @johnathanoddie @isry98

OpenStudy (anonymous):

u called?

OpenStudy (theweeknd16):

yeah.

OpenStudy (ccchristian1):

whats the answer

OpenStudy (ccchristian1):

@too_smart143

OpenStudy (ccchristian1):

a or D @too_smart143

OpenStudy (anonymous):

i cant give u a direct answer sorry i got reported for it

OpenStudy (anonymous):

which one do u think it is

OpenStudy (ccchristian1):

i think its D

OpenStudy (ccchristian1):

am i right or wrong

OpenStudy (ccchristian1):

@too_smart143

OpenStudy (anonymous):

ig